# A minimal coalition logic

### v2

@@ Abstract
Coalition Logic is a central logic in logical studies of strategic reasoning, whose models are concurrent game models. In this paper, first, we systematically discuss three assumptions of concurrent game models and argue that they are too strong. The first is seriality; that is, every coalition always has an available joint action. The second is the independence of agents; that is, the merge of two available joint actions of two disjoint coalitions is always an available joint action of the union of the two coalitions. The third is determinism; that is, all available joint actions of the grand coalition always have a unique outcome. Second, we present a coalition logic based on general concurrent game models which do not have the three assumptions and show its completeness. This logic seems minimal for reasoning about coalitional powers.
@@ 1 Introduction
Coalition Logic $\mathsf{CL}$ was proposed by Pauly [8, 9]. Many logics about strategic reasoning are extensions of $\mathsf{CL}$. For example, Alternating-time Temporal Logic $\mathsf{ATL}$ ([1]) is a temporal extension of $\mathsf{CL}$; Strategy Logic $\mathsf{SL}$ [6] is an extension of $\mathsf{CL}$, whose language has quantifiers for and names of strategies. We refer to [2], [10], and [13] for overviews of the area.
@@
The language of $\mathsf{CL}$ is a modal language with the featured operator $\langle[\mathrm{A}]\rangle\phi$, indicating some available joint action of the coalition $\mathrm{A}$ ensures $\phi$. Its models are concurrent game models. Roughly, in a concurrent game model: there are some states; there are some agents who form coalitions; at any state, every coalition has some available joint actions; every joint action of a coalition has some possible outcome states. The formula $\langle[\mathrm{A}]\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{A}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action in the model.
@@
Concurrent game models have some assumptions and three of them are too strong. The first one is seriality: every coalition always has an available joint action. The second one is the independence of agents: the merge of two available joint actions of two disjoint coalitions is always an available joint action of the union of the two coalitions. The third one is determinism: every available joint action of the grand coalition has a unique outcome.
In the research line of STIT (see to it that) logic, Sergot has explicitly argued against the independence of agents in [11], and has explicitly argued that there are non-deterministic scenarios in [12]. Boudou and Lorini [3] introduced new semantics for temporal STIT logic, where non-deterministic concurrent game models are used. In the research line of coalition logic, Jiang and Naumov [5] proposed the use of non-serial and non-deterministic models.  [^2] In their models, Jiang and Naumov [5] allowed an action profile to have no outcome state or more than one outcome state at a state; they also allowed dead ends, that is, states where no action profile has an outcome state. However, they did not explicitly define available joint actions. There are two ways to understand this. One is that there is no such notion of available joint actions. Then it makes no sense to say their models are serial/deterministic or not, as seriality and determinism are about available joint actions. Another way is that they implicitly treated all joint actions, including those without any outcome state, as available. Then, their models might be non-deterministic but are serial.
@@
In this paper, first, we systematically discuss the three assumptions of concurrent game models and argue that they are too strong. Second, based on general concurrent game models, which do not have the three assumptions, we present a Minimal Coalition Logic $\mathsf{MCL}$, compare it to $\mathsf{CL}$ in detail, and show its completeness.
@@
The rest of the paper is structured as follows. In Section 2, we present some general settings of coalition logics. In Section 3, we give $\mathsf{CL}$ in detail. In Section 4, we discuss the three assumptions and show that they do not hold in all situations. In Section 5, we present $\mathsf{MCL}$, argue that $\mathsf{MCL}$ is a minimal coalition logic for reasoning about coalitional powers, and discuss whether $\mathsf{CL}$ can simulate $\mathsf{MCL}$. We show the completeness of $\mathsf{MCL}$ in Section 6. We point out some further work in Section 7. We put the proofs for some results in the Appendix.
@@ 2.1 Language
$\phi::=\top\mid p\mid\neg\phi\mid(\phi\wedge\phi)\mid\langle[\mathrm{A}]\rangle\phi$
The formula $\langle[\mathrm{A}]\rangle\phi$ indicates that some available joint action of $\mathrm{A}$ ensures $\phi$.
@@
  • The propositional connectives ,$\bot$,$\lor$,$\rightarrow$ , ∨ , →, and $\leftrightarrow$ are defined as usual.
  • Define the dual $[\langle\mathrm{A}\rangle]\phi$ of $\langle[\mathrm{A}]\rangle\phi$ as $\neg\langle[\mathrm{A}]\rangle\neg\phi$, indicating that every available joint action of $\mathrm{A}$ enables $\phi$.
  • $\Box\phi$, defined as $\langle[\emptyset]\rangle\top\rightarrow\langle[\emptyset]\rangle\phi$, means that $\phi$ will necessarily be true.
  • $\Diamond\phi$, defined as $\langle[\emptyset]\rangle\top\land[\langle\emptyset\rangle]\phi$, means that $\phi$ will possibly be true.
@@ 2.2 Abstract multi-agent action models
    As joint actions are treated as unary functions, which are sets of binary tuples, set-theoretical notions such as union and intersection are applicable to them.
@@
We understand available joint actions as follows: a joint action $\sigma_{\mathrm{A}}$ of a coalition $\mathrm{A}$ is available at a state $s$ if and only if $\sigma_{\mathrm{A}}$ has an outcome state at $s$. We say that a joint action of a coalition is conditionally available if whether it is available is dependent on other agents’ actions simultaneously.
@@ 2.3 Semantics
$\mathtt{M},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
$\mathtt{M},s\Vdash\neg\phi$Leftrightarrow not $\mathtt{M},s\Vdash\phi$
$\mathtt{M},s\Vdash\phi\land\psi$Leftrightarrow $\mathtt{M},s\Vdash\phi$ and $\mathtt{M},s\Vdash\psi$
$\mathtt{M},s\Vdash\langle[\mathrm{A}]\rangle\phi$Leftrightarrow there is $\sigma_{\mathrm{A}}\in\mathtt{av}_{\mathrm{A}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{A}}(s,\sigma_{\mathrm{A}})$, $\mathtt{M},t\Vdash\phi$
@@
$\mathtt{M},s\Vdash[\langle\mathrm{A}\rangle]\phi$Leftrightarrow for all $\sigma_{\mathrm{A}}\in\mathtt{av}_{\mathrm{A}}(s)$, there is $t\in\mathsf{out}_{\mathrm{A}}(s,\sigma_{\mathrm{A}})$ such that $\mathtt{M},t\Vdash\phi$
$\mathtt{M},s\Vdash\Box\phi$Leftrightarrow if $\emptyset\in\mathtt{av}_{\emptyset}(s)$, then for all $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$, $\mathtt{M},t\Vdash\phi$
$\mathtt{M},s\Vdash\Diamond\phi$Leftrightarrow $\emptyset\in\mathtt{av}_{\emptyset}(s)$ and there is $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$ such that $\mathtt{M},t\Vdash\phi$
@@ 3.1 Concurrent game models
Let $\{\mathrm{A}_{i}\mid i\in I\}$ be a family of pairwise disjoint coalitions for some (possibly empty) index set $I$, and $\Lambda=\{\Sigma_{\mathrm{A}_{i}}\mid i\in I\text{ and }\Sigma_{\mathrm{A}_{i}}\text{is a set of joint actions of }\mathrm{A}_{i}\}$.
@@
$\bigoplus\Lambda=\{\sigma:\bigcup_{i\in I}\mathrm{A}_{i}\to\mathtt{AC}\mid\text{there is a choice function $f$ of }\Lambda\text{ such that }\sigma=\bigcup_{i\in I}f(i)\},$
@@
The proof for this result is put to the Appendix.

### v3

@@ Abstract
Coalition Logic is an important logic in logical studies of strategic reasoning, whose models are concurrent game models. In this paper, first, we systematically discuss three assumptions of concurrent game models and argue that they are too strong. The first is seriality; that is, every coalition always has an available joint action. The second is the independence of agents; that is, the merge of two available joint actions of two disjoint coalitions is always an available joint action of the union of the two coalitions. The third is determinism; that is, all available joint actions of the grand coalition always have a unique outcome. Second, we present a coalition logic based on general concurrent game models that do not have the three assumptions and show its completeness. This logic seems minimal for reasoning about coalitional powers.
@@ 1 Introduction
Coalition Logic $\mathsf{CL}$
@@
Coalition Logic $\mathsf{CL}$ is a logic for reasoning about coalitional powers, proposed by Pauly [12, 13]. The language of $\mathsf{CL}$ is a modal language with the featured operator $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$, indicating some available joint action of the coalition $\mathrm{A}$ ensures $\phi$. Its models are concurrent game models. Roughly, in a concurrent game model: there are some states; there are some agents who form coalitions; at any state, every coalition has some available joint actions; every joint action of a coalition has some possible outcome states. The formula $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{A}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action.
Coalition Logic $\mathsf{CL}$ is an important logic for strategic reasoning. Many logics in this field are extensions of $\mathsf{CL}$. For example, Alternating-time Temporal Logic $\mathsf{ATL}$ ([1]) is a temporal extension of $\mathsf{CL}$, and strategy Logic $\mathsf{SL}$ [9] is an extension of $\mathsf{CL}$, whose language has quantifiers for and names of strategies. We refer to [3], [14] and [18] for overviews of the area.
Our work
@@
Concurrent game models rest on several assumptions. In this paper, we consider three of them. The first one is seriality: coalitions always have available joint actions. The second one is the independence of agents: the merge of two available joint actions of two disjoint coalitions is always an available joint action of the union of the two coalitions. The third one is determinism: available joint actions of the grand coalition always have a unique outcome.
@@
In this work, we first systematically discuss the three assumptions and argue that they are too strong. Second, present a Minimal Coalition Logic $\mathsf{MCL}$ based on general concurrent game models which do not have the three assumptions, compare $\mathsf{MCL}$ to $\mathsf{CL}$ in detail, and show its completeness.
The logic $\mathsf{MCL}$ is minimal in the following sense: general concurrent game models cannot be more “general”, given that they are for coalitional powers.
Related work
There has been some work in the agency literature that drops some of the three assumptions.
Jiang and Naumov [8] presented a logic for reasoning about knowledge and strategies in multi-agent systems, whose framework is in the tradition of coalition logic. In their models, Jiang and Naumov allowed an action profile to have no outcome state or more than one outcome state. They also allowed dead ends, that is, states where no action profile has an outcome state  [^2] In [8], Jiang and Naumov did not explicitly define available joint actions. There are two ways to understand this. One is that there is no such notion of available joint actions. Then it makes no sense to say their models are serial/deterministic or not, as seriality and determinism are about available joint actions. Another way is that they implicitly treated all joint actions, including those without any outcome state, as available. Then, their models are non-deterministic but are serial. . They did this to model termination of games.
The independence of agents commonly holds in STIT (see to it that) logics. Sergot [16] found no convincing justification for this principle in the literature of STIT logics. He offered examples showing that the principle fails. Based on an extension of Propositional Dynamic Logic with multi-agents, Royakkers and Hughes [15] formalized three notions of responsibilities. They showed that the independence of agents does not hold generally and argued that it is also inappropriate for one notion of responsibility, that is, accountability. They use models where the independence of agents fails. In addition, determinism does not hold in their models either. They allowed this for the reason that agents may perform actions with unknown outcomes.
Boudou and Lorini [4] introduced new semantics for a temporal STIT logic, utilizing non-deterministic concurrent game models. Sergot [17] investigated how causal responsibility can be treated in STIT logics, where he explicitly argued that there are non-deterministic scenarios. To deal with outcome uncertainty of agents’ actions, some work, such as [5] and [10], has introduced probability to concurrent game models. In these work, available joint actions of the grand coalition at a state may have different outcome states with different probabilities.
Our brief reasons to drop the three assumptions
Many artificial settings (such as games) have terminal states, where agents do not have available actions. To model terminal states of these artificial settings, we need to give up seriality.
Agents’ actions often interact with each other, and whether an agent can perform an action is often conditional on other agents’ actions at the same time. In these situations, the independence of agents fails. To model agents’ conditional abilities, we need to drop the assumption of agent independence.
There are many situations where some joint actions of all behaving agents have more than one outcome state. Additionally, in many situations, we just consider some but not all agents who form the grand coalition. In these situations, determinism does not hold.
Structures of the paper
@@
The rest of the paper is structured as follows. In Section 2, we present some general settings of coalition logics. In Section 3, we give $\mathsf{CL}$ in detail. In Section 4, we discuss the three assumptions and show that they do not hold in all situations. In Section 5, we present $\mathsf{MCL}$, argue that $\mathsf{MCL}$ is a minimal coalition logic for reasoning about coalitional powers, and discuss whether $\mathsf{CL}$ can simulate $\mathsf{MCL}$. We show the completeness of $\mathsf{MCL}$ in Section 6. We point out some further work in Section 7.
@@ 2.1 Language
$\phi::=\top\mid p\mid\neg\phi\mid(\phi\wedge\phi)\mid\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$
The formula $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$ indicates that some available joint action of $\mathrm{A}$ ensures $\phi$. Here, the notation of $\langle\hskip-2.0pt[\cdot]\hskip-2.0pt\rangle\phi$ differs from its notation in the literature. We do this to indicate the iteration of quantifiers in its meaning, which will be seen below.
@@
  • The propositional connectives $\bot,\lor,\rightarrow$, and $\leftrightarrow$ are defined as usual.
  • Define the dual $\hskip-2.5pt\langle\mathrm{A}\rangle\hskip-2.5pt]\phi$ of $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$ as $\neg\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\neg\phi$, indicating that every available joint action of $\mathrm{A}$ enables $\phi$.
[…]
@@ 2.2 Abstract multi-agent action models
    As joint actions are treated as unary functions, which are sets of binary tuples, set-theoretical notions such as union and intersection apply to them.
@@
We understand available joint actions as follows: a joint action $\sigma_{\mathrm{A}}$ of a coalition $\mathrm{A}$ is available at a state $s$ if and only if $\sigma_{\mathrm{A}}$ has an outcome state at $s$. We say that a joint action of a coalition is conditionally available if its availability depends on the simultaneous actions of other agents.
@@ 2.3 Semantics
$\mathtt{M},s\Vdash p$ $\Leftrightarrow$ $p\in\mathsf{label}(s)$
$\mathtt{M},s\Vdash\neg\phi$ $\Leftrightarrow$ not $\mathtt{M},s\Vdash\phi$
$\mathtt{M},s\Vdash\phi\land\psi$ $\Leftrightarrow$ $\mathtt{M},s\Vdash\phi$ and $\mathtt{M},s\Vdash\psi$
$\mathtt{M},s\Vdash\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$ $\Leftrightarrow$ there is $\sigma_{\mathrm{A}}\in\mathtt{av}_{\mathrm{A}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{A}}(s,\sigma_{\mathrm{A}})$, $\mathtt{M},t\Vdash\phi$
@@
$\mathtt{M},s\Vdash[\hskip-2.5pt\langle\mathrm{A}\rangle\hskip-2.5pt]\phi$ $\Leftrightarrow$ for all $\sigma_{\mathrm{A}}\in\mathtt{av}_{\mathrm{A}}(s)$, there is $t\in\mathsf{out}_{\mathrm{A}}(s,\sigma_{\mathrm{A}})$ such that $\mathtt{M},t\Vdash\phi$
$\mathtt{M},s\Vdash\Box\phi$ $\Leftrightarrow$ if $\emptyset\in\mathtt{av}_{\emptyset}(s)$, then for all $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$, $\mathtt{M},t\Vdash\phi$
$\mathtt{M},s\Vdash\Diamond\phi$ $\Leftrightarrow$ $\emptyset\in\mathtt{av}_{\emptyset}(s)$ and there is $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$ such that $\mathtt{M},t\Vdash\phi$
@@ 3.1 Concurrent game models
[…]
@@
[…]
@@
    This constraint indicates that concurrent game models are serial: every coalition always has an available joint action.
@@
    This constraint indicates that concurrent game models are deterministic: every available joint action of the grand coalition has a unique outcome.
@@
Proof.
Let $\mathrm{A}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$.
Assume $\mathrm{A}=\emptyset$.
Assume $\mathrm{A}=\{a_{1},\dots,a_{n}\}$.